OpenStudy (anonymous):

If (x+1)(x+2)(x+3) = 720 then A) (X+3) B) 10 which one is greater Option A or Option B...?

OpenStudy (anonymous):

option B for whatever you plug in for x, -1 -2 or -3

OpenStudy (anonymous):

B is greater

OpenStudy (anonymous):

Sorry 3rd option is also there C) two are equal..

OpenStudy (anonymous):

Hear it is given that two options are equal that is x+3=10 i dont know how it is....?

OpenStudy (anonymous):

@Callisto (Moderator)

OpenStudy (anonymous):

@Zarkon

OpenStudy (kc_kennylau):

Well the question is wrong itself, so...

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (tyteen4a03):

(x+1)(x+2)(x+3) - 720 = 0 Expand and simplify, then try to find a common factor and factorize.

OpenStudy (anonymous):

yes that is x^3+6x^2+11x-714=0 then what to do...?

OpenStudy (anonymous):

@tyteen4a03

OpenStudy (tyteen4a03):

Factorize to get rid of the x^3 term. I am not a factorizer so I use W|A, but you should get something like this: (x - stuff)(x^2 + stuff)

Directrix (directrix):

@chetan552 Is there an option C stating that Option A = Option B?

OpenStudy (agent0smith):

(x+1)(x+2)(x+3) = 720 then A) (X+3) B) 10 Why not just let x+3 = 10, then x+2=9 and x+1=8 10*9*8 = 720

OpenStudy (anonymous):

Yes option c is also there Two are equal...@Directrix

OpenStudy (anonymous):

It is not eassy to guess forother question...@agent0smith

OpenStudy (agent0smith):

"Why not just let x+3 = 10, then x+2=9 and x+1=8 10*9*8 = 720" That's a way to check whether 10 is greater x+3 or not. If the 10*9*8 turned out to be less than 720, then it would mean x+3 must have to be greater than 10.

OpenStudy (anonymous):

HOw did you get the answer...@Directrix

OpenStudy (agent0smith):

It's not a guess @chetan552 it's reasoning.

OpenStudy (agent0smith):

(x+1)(x+2)(x+3) = 1024 then A) (X+3) B) 16 which is greater? Find out. Set x+3 = 16, then x+2 = 15, x+1 = 14 16*15*14 = 3360 so 16 is much larger than x+3

OpenStudy (anonymous):

your answer is correct @agent0smith but how did you say that...

OpenStudy (agent0smith):

I got it the same way I got the one above. Checking which is bigger.

OpenStudy (anonymous):

Thank you i got it.....@agent0smith

OpenStudy (agent0smith):

(x+1)(x+3)(x+5) = 525 then A) (X+5) B) 10 which is greater? Find out the same way as above... 10*8*6 = 480. So x+5 must be bigger than 10.

OpenStudy (anonymous):

your answer is correct @agent0smith but how did you say that...

OpenStudy (agent0smith):

The key is to understand that the brackets must be factors of 720. If (x+1)(x+2) = 12 Then the numbers in parentheses must be factors of 12, right? Like 1*12 = 12 or 2*6 = 12 or 3*4 = 12

OpenStudy (agent0smith):

They could also be decimals (usually in questions like the SAT they'll state that they're integers), but either way, they're two numbers that multiply to 12

OpenStudy (anonymous):

I got it Thank you.....@agent0smith